Eric Gunther

Cambridge, MA USA

The Vibravibe VL-12 (2005)

Vibrotactile lounge chair: plywood, canvas, lycra velvet, low-frequency vibrotactile transducers. Composition: music and vibrotactile compositions, control software.

Vibrotactile stimulation is often pigeonholed into the realms of pleasure and relaxation. Looking at vibration as a compositional medium - with the vibrations themselves as aesthetic artifacts - we begin to see how many of the dynamic, rhythmic, and harmonic structures of sonic vibrations in music can be naturally extended to palpable vibrations. Space and motion on the vast surface of the skin are essential parameters and the body becomes the stage for a dance that is felt rather than seen. Thanks to Jeff Lieberman, Justin Manor, Chris Parlato, John Rothenberg, and Small Design Firm for their help with this project.

Steve Hollinger

Boston, MA USA

www.sjh.com

House of Style (2004)

mixed media sculpture

At intervals of approximately two minutes, House of Style will randomly select and present a fashion suggestion from an array of styles. In its normal environment, House of Style is placed in a window and operates during the day, drawing energy from sunlight. If you are disatisfied with the current style on display, please feel free to return in a few minutes.

jackbackrack

Cambridge, MA USA

www.jbot.org

Protochoice (2005)

Modular light: 1x16” PCBs, microcontrollers, LEDs, and USB cables and connectors.

Protochoice is a modular lighting prototype which computes a lifetime of chance decisions. It is comprised of 26 1x16” printed circuit boards with two hot pluggable USB connections on each end providing both structural and electical glue. A collection of protochoice boards permit the construction of a wide range of wireframe sculptures. Protochoice is the first in a series of modular electronic sculptures by the artist. Many thanks to Mark Tobenkin, Jeff Lieberman, Kevin McCormick, James Patten, Hayes Raffle, Fran Trainor, Brian Knep, Dan Paluska, Mindy Zarem, and the Collision Collective. Inspired by Eric Saund’s Markov cube and XTC’s “Complicated Game” song (i.e., should I part my hair on the left or right?). Partially funded by the MIT Council for the Arts.

Jeff Lieberman, Josh Lifton, David Merrill, Hayes Raffle

Cambridge, MA USA

Applause (2005)

Mini movie theater, archival film footage, electronics

An interactive video installation draws on classic film archives to explore contemporary issues surrounding images, advertising, and group participation. Please assemble a small group and applaud to begin the show. Extra thanks: Jack B., Dan P., Brian K., Joshua G., Sajid S., Mark F., and Paloma.

Michael Mittelman

Boston, MA USA

Performance 0405 (2005)

Sound

A mixture of live, delayed, and pre-recorded sound creates a quadraphonic soundscape that is individual to each listener and each moment.

Gretchen Skogerson, Garth Zeglin

Somerville, MA USA and Pittsburgh, PA USA

www.publicsecrets.net

PS (2005)

mixed media (mirror, electronics)

Lean in close.

Elisabeth Sylvan

Somerville, MA USA

http://www.media.mit.edu/~sylvan

Bodies of Light (2005)

Light and motors

How can life be drawn from a machine? In “Bodies of Light” servo-controlled white lights push against white fabric. They writhe and contort, creating kaleidoscopic patterns.

William Tremblay, Rob Gonsalves

Allston, MA USA

http://www.williamtremblay.com/art/janken/

Janken (2005)

Interactive computer graphics, projection, video camera, plexiglas, sensors, rock, paper and scissors

Janken is the Japanese word for the common game known in the US as “Rock-Paper-Scissors”. Roughly translated, it means “hand game” but also implies fist, strike and gambling.

Noah Vawter

Cambridge, MA USA

web.media.mit.edu/~nvawter

1-Bit Love (2005)

Musical altar with big, comfy knobs

We all have fallen under the spell of 1-bit waveforms. With some increases in music technology, we have forgotten about them briefly. However, the memories rest in our hearts...along with a desire to hear their many variations which no one else has explored.

Orit Zuckerman, Sajid Sadi

Cambridge, MA USA

Influence (2005)

Digital portraits, projection, glass

How individuals are affected by the behavior of people around them.